OpenStudy (anonymous):

3x + 13 (3x + 13) = 19, solve for x

hero (hero):

3x + 39x + 169 = 19 42x + 169 = 19 42x = 19 - 169 42x = -150 x = -150/42 x = -75/21 x = -25/7